# Hyperspectral Image Generation with Unmixing Guided Diffusion Model

### v1

@@ Abstract
Recently, hyperspectral image generation has received increasing attention, but existing generative models rely on conditional generation schemes, which limits the diversity of generated images. Diffusion models are popular for their ability to generate high-quality samples, but adapting these models from RGB to hyperspectral data presents the challenge of high dimensionality and physical constraints. To address these challenges, we propose a novel diffusion model guided by hyperspectral unmixing. Our model comprises two key modules: an unmixing autoencoder module and an abundance diffusion module. The unmixing autoencoder module leverages unmixing guidance to shift the generative task from the image space to the low-dimensional abundance space, significantly reducing computational complexity while preserving high fidelity. The abundance diffusion module generates samples that satisfy the constraints of non-negativity and unity, ensuring the physical consistency of the reconstructed HSIs. Additionally, we introduce two evaluation metrics tailored to hyperspectral data. Empirical results, evaluated using both traditional metrics and our proposed metrics, indicate that our model is capable of generating high-quality and diverse hyperspectral images, offering an advancement in hyperspectral data generation.
@@ I Introduction
Hyperspectral image (HSI) generation represents a critical need in remote sensing. HSI is significant in various applications [1, 2], but the acquisition of high-fidelity HSIs is constrained by limitations in sensor technology [3]. These challenges lead to the scarcity of large-scale, high-quality HSI data, hindering the advancements in hyperspectral research [4]. Consequently, synthesizing diverse HSIs has emerged as a pivotal yet underexplored task.
However, existing methods cannot expand data due to their conditional generation frameworks [5]. These approaches typically employ conditional constraints, such as RGB images or segmentation maps, to guide the distribution of generated hyperspectral images [6, 7]. While effective for spectral super-resolution, these methods fall short of generative modeling, as they prioritize spectral restoration over the comprehensive simulation of the underlying HSI distribution. Although conditional frameworks ensure high-fidelity outputs, they inherently limit the diversity of generated images.
@@
To address these challenges, we integrate hyperspectral unmixing into the generative framework. We project HSIs into the abundance space, which is low-dimensional and governed by the explicit constraints, namely non-negativity and unity [14]. By operating in the abundance space, a generative model can effectively synthesize hyperspectral data while maintaining physical consistency.
To implement the proposed framework, we introduce the Hyperspectral Unmixing Diffusion Model (HUD), which consists of two key components: an Unmixing AutoEncoder (UAE) pair and an Abundance Diffusion Module (ADM). To address the challenge of high dimensionality, we design the UAE module. UAE is initialized using a hyperspectral unmixing algorithm, where abundances are derived either through rigorously constrained optimization equations or a linear approximation. Reconstruction is performed directly using the endmembers, ensuring efficient dimensionality reduction while preserving essential spectral information. To address the challenge of physical constraint, we construct the ADM module. The diffusion model iteratively generates samples on the abundance space. However, traditional diffusion models rely on Gaussian distributions to describe latent features, which inherently violate the unity and non-negativity constraints of the abundance space. To address this limitation, we project the abundance space into an unconstrained domain, apply Gaussian-based sampling, and then map the generated samples back to the abundance space. This approach enables the generation of physically consistent and high-fidelity hyperspectral data while maintaining computational efficiency. Moreover, we introduce two novel metrics tailored for HSI generation: point fidelity and block diversity. Experimental results indicate that HUD consistently outperforms existing models on both proposed metrics and traditional metrics.
@@ II Related Work
In this section, we review diffusion models and existing methods for HSI generation. We also introduce the hyperspectral unmixing algorithm.
@@ II-A Diffusion Models
Denoising Diffusion Probabilistic Models (DDPM) [15] have recently emerged as a hot topic in the field of generative networks. DDPM constructs a diffusion random process that gradually changes the image distribution into the standard Gaussian noise, and a denoising reverse process that changes the noise into an image. The denoising process is a Markov chain, and the transition probability is approximated by a U-Net [16].
Latent Diffusion Model [17], also known as Stable Diffusion, is an improvement of DDPM. The dimension of the intermediate variables during the denoising process of DDPM is constant. As a result, DDPM is faced with the problem of high computational expense and low image resolution. Stable diffusion encodes the original image into low-dimensional latent space and constructs a DDPM on it. The autoencoder is pretrained on ImageNet or other large-scale RGB datasets.
@@ II-B HSI Generation
Recently, several studies have explored the use of generative models for HSI generation, including GAN, VAE, and Stable Diffusion [25, 26, 27, 28, 29]. However, most of these approaches still rely on conditional images such as RGB or multi-spectral images during the generation process. This reliance limits the diversity of the generated hyperspectral images, as the spatial distribution is largely determined by the conditional input.
@@ III Method
Figure 1: Components of HUD. HUD is composed of 2 modules: the unmixing autoencoder and the abundance diffusion model.
In this section, we show the detailed design of the proposed model, including the unmixing autoencoder module and the abundance diffusion module. We also introduce the evaluation metrics for hyperspectral image generation. The overall structure of HUD is shown in Figure 1.
III-A Unmixing Autoencoder Module
The unmixing autoencoder can use any algorithm to extract endmember, and in this paper we use VCA for example. Generally, abundance is typically subject to two key constraints: non-negativity and unity. Specifically, the abundance values for each pixel must be non-negative and sum to one, reflecting the physical interpretation that each pixel is composed of a mixture of endmembers in varying proportions. Therefore, given $Y$ and $A$, $X$ is usually solved through a constrained optimization problem of the following form:
$\min||Y-AX||^{2}\text{\, s.t.\, }X>0\text{\, and\, }\sum_{i}{X_{(i,:,:)}=\mathbbm{1}_{h\times w}}$ (2)
@@ III-A Unmixing Autoencoder Module
III-B Abundance Diffusion Module
@@
where $Z$ represents the projected feature in an unconstrained space, and $\hat{X}$ is the reconstructed abundance map that satisfies the original constraints. This projection is designed to be lossy, with the difference between $X$ and $\hat{X}$ being negligible:
@@
For simplification, we assume that $\Sigma_{\theta}(Z_{t},t)=\sigma^{2}_{t}\mathbf{I}$, and use reparameterizing $Z_{t}(Z_{0},\epsilon)=\sqrt{\bar{\alpha}_{t}}Z_{0}=\sqrt{1-\bar{\alpha}_{t}}\epsilon$ for $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, the middle term in Equation 12 can be written as:
@@
$L:=\mathbf{E}_{t,Z_{0},\epsilon}\left[||\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}Z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)||^{2}\right]$ (15)
@@
After training, we sample from standard Gaussian distribution to get $Z_{T}$ and follow the denoising process step by step. The training and sampling procedure are shown in Algorithms 1 and 2.
@@
     make backward propagation by $\triangledown_{\theta}||\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}Z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)||^{2}$
@@
III-C Point Fidelity and Block Diversity
@@
$F_{p}=\frac{1}{|\hat{Y}|}\sum_{\hat{y}\in\hat{Y}}\max_{y\in Y}\left[\frac{\hat{y}}{||\hat{y}||}*\frac{y}{||y||}\right]$ (16)
where $Y$ is the real HSI, $\hat{Y}$ is the generated HSI, $y$ and $\hat{y}$ are pixels, and $|\hat{Y}|$ is the amount of pixels. In contrast, block diversity reflects whether the overall distribution of the generated HSI is a direct copy of the original HSI, or if there are differences in the distribution during the generation process. To be specific:
$D_{b}=\frac{1}{N_{b}}\sum_{\hat{y}_{b}\subset\hat{Y}}\frac{1}{|\hat{y}_{b}|}\max_{y_{b}\in X}\left[\frac{\hat{y}_{b}}{||\hat{y}_{b}||}*\frac{y_{b}}{||y_{b}||}\right]$ (17)
@@ IV Experiments
In this section, we empirically demonstrate the superiority of HUD. We first introduce the datasets used in the experiments, the comparison models, and the selection of hyperparameters. Then, we list images generated by different models. Finally, we analyze the performance of different models using quantitative metrics.
@@ IV-A Experiments Setup
We use Indianpines, KSC, Pavia, PaviaU, and Salinas to showcase the generation results. (1) Indianpines contains an image scene with spatial size $145\times 145$ and 220 bands, covering a wavelength range of $[104-108],[150-163],220$ nm, having 16 classes. The Band 46,17,11 from the data are chosen for pseudo-color visualization. (2) KSC contains an image scene with spatial size $512\times 614$ and 176 bands, covering a wavelength range of 400 - 2500 nm, having 13 classes. The Band 28,9,10 from the data are chosen for pseudo-color visualization. (3) Pavia contains an image scene with spatial size $1093\times 715$ and 102 bands, having 9 classes. The Band 46,27,10 from the data are chosen for pseudo-color visualization. (4) PaviaU contains an image scene with spatial size $610\times 340$ and 103 bands, having 9 classes. The Band 46,27,10 from the data are chosen for pseudo-color visualization. (5) Salinas contains an image scene with spatial size $512\times 217$ and 204 bands, covering a wavelength range of $[104-108],[150-163],220$ nm, having 16 classes. The Band 36,17,11 from the data are chosen for pseudo-color visualization.
HUD will be compared with the following models: VAE, GAN [40], and MPRNet [41]. As there are currently no unconditional hyperspectral image generation algorithms, we adapted the classic models VAE and GAN from RGB generation tasks. Their first convolutional layer, usually a 1x1 conv layer, is resized according to the dimensions of the dataset. The other structures remain unchanged. We also evaluate the generation performance of MPRNet, a state-of-the-art hyperspectral image generation model, but it requires RGB images as conditions. Following the original paper, we directly input the pseudo RGB images to MPRNet.
The UAE of HUD is initialized separately for each dataset, and the dimensionality of $Z$ is determined by the number of categories in the dataset. The UAE is pretrained and does not participate in the training of the diffusion model. The time steps of the diffusion model are set to 1000. Except for the downsampling layer, the structures and hyperparameters of VAE, GAN, and MPRNet are set according to the original papers.
All models are trained on a 4090 GPU. Although we designed the UAE to reduce the computational load of HUD, it still cannot generate full-sized hyperspectral images. Additionally, the comparison algorithms also fail to generate complete images. Therefore, during training, we randomly crop 32x32 overlapping sub-images from the original image as training samples, and the generation process also produces 32x32 images.
@@ IV-B1 Results on Pseudo-Color Images
The pseudo-color visualization of 5 datasets is shown in Figure 3. The first column shows the pseudo-color visualization of the real HSI, and the generated images from different methods are listed on the right side. The pseudo-color visualization can reflect the spatial distribution of HSIs. Specifically, the generated images from GAN have grids overlay on the surface, resembling a mosaic. This problem is probability caused by the downsampling layers in generators and discriminators, which overlook some specific areas because of the pooling or striding operation. VAE directly estimates the distribution of original images, but the scheme limits the performance of the model. Consequently, the generated images from VAE are blurry and lack details. As a spectral super-resolution model, MPRNet generates images with high quality, but the spatial distribution is limited to the original image. HUD also generates high-quality images, and the spatial distribution is more diverse than MPRNet. For example, there are some new roads and buildings in the generated images of HUD, which are not present in the original images.
IV-B2 Results on Spectral Curves
@@
The spectral curve distributions of the real HSI and the generated HSI are shown in Figure 4. We stack all the spectral curves of an HSI to show the spectral distributions. The boundary shows whether the generated spectral curves are consistent with the real HSI, and the density (color) infers the distribution of endmembers. Generally, GAN, MPRNet and HUD generate spectral curves that are more consistent with the real HSI than VAE. GAN tends to generate spectral curves that are frequently seen in the real HSI to confuse the discriminator by high-quality samples, so the distribution of generated spectral curves are concentrated. VAE only estimates the main characteristics of the real HSI due to the deficient structure. MPRNet generates high-quality spectral curves, but the distribution is limited to the real HSI, so the generated images lack diversity. HUD generates high-quality spectral curves that follows the overall shape and distribution of the real HSI, while there are also different distributions in the detail.
@@ IV-C1 Results on Fidelity and Diversity
$F_{p}$uparrow $D_{b}$downarrow $D_{b}/F_{p}$downarrow
Dataset GAN VAE MPRNet HUD GAN VAE MPRNet HUD GAN VAE MPRNet HUD
Indianpines 0.998 0.914 0.996 0.999 0.981 0.914 0.995 0.986 0.983 1.000 0.999 0.987
KSC 0.998 0.978 0.983 0.999 0.976 0.978 0.983 0.990 0.988 1.000 1.000 0.991
Pavia 0.996 0.941 0.999 0.999 0.949 0.936 0.942 0.912 0.953 0.995 0.943 0.913
PaviaU 0.996 0.941 0.997 0.999 0.948 0.934 0.984 0.919 0.952 0.992 0.987 0.920
Salinas 0.995 0.998 0.998 0.999 0.984 0.982 0.988 0.981 0.989 0.984 0.990 0.982
Avg 0.994 0.954 0.995 0.999 0.968 0.949 0.978 0.958 0.969 0.994 0.984 0.959
The quantitative results are shown in Tables II, III and I. According to Table II, all models achieve similar results while HUD performs slightly better than the other models. According to Table III, the results are unstable across different datasets and models. IS and FID are originally proposed for RGB images, which requires a generally pretrained InceptionV3. However, the hyperspectral images have significant differences in distribution, so the classification networks should be trained on each dataset. As a result, the metrics are unstable and questionable. Regardless, HUD and GAN achieve the top results in the IS and FID comparison.
@@
Metric Dataset GAN VAE MPRNet HUD
IS $\uparrow$ Indianpines 1.08 1.02 1.04 1.07
KSC 1.00 1.00 1.00 1.00
Pavia 1.01 1.01 1.04 1.01
PaviaU 1.03 1.03 1.04 1.05
Salinas 1.05 1.04 1.00 1.05
Avg 1.03 1.02 1.03 1.04
@@
Metric Dataset GAN VAE MPRNet HUD
FID $\downarrow$ Indianpines 8.03 240.91 20.98 4.71
KSC 10.87 5.45 4.76 6.69
Pavia 4.81 13.82 2.37 4.79
PaviaU 4.92 9.13 3.13 2.81
Salinas 15.03 21.33 29.09 23.12
Avg 8.73 58.13 12.07 8.42
As shown in Table I, HUD achieves the highest point fidelity, indicating that its generated images better preserve the spectral content of real images. MPRNet, a spectral super-resolution model, ranks second in point fidelity. In terms of block diversity, VAE generates the most diverse images, but its poor image quality results in lower point fidelity. HUD, however, achieves the best block diversity after VAE, demonstrating richer spatial distributions. MPRNet ranks lowest in block diversity, suggesting its generated images are overly constrained by the original spatial patterns. Combining both point and block diversity, HUD delivers the best overall performance, followed by GAN. Despite VAE excelling in block diversity, its overall performance remains the weakest.
IV-C2 Results on Computational Complexity
Due to the high dimensionality of hyperspectral images, we designed an autoencoder with unmixing guidance to reduce the computational load of the diffusion model. Despite the efforts, we were only able to generate 32x32 hyperspectral images in our experiments. This section compares the computational costs of different models for generating Indianpines images, including a hypothetical estimate for DDPM. The results are summarized in Table IV.
TABLE IV: Computational Complexity.
Size Metric GAN VAE MPRNet HUD DDPM
32 Params GB 0.49 0.49 0.56 1.15 1.43
Memory GB 1.67 1.81 1.97 3.35 72.22
FLOPs $10^{11}$ 0.19 0.20 0.22 46.75 1038.21
256 Params GB 0.49 0.49 0.56 1.15 1.43
Memory GB 9.93 11.05 12.43 18.75 567.75
FLOPs $10^{11}$ 1.52 1.60 1.76 374.00 8305.68
Diffusion models require more parameters (about double of the others) due to the full U-Net structure used for predicting noisy samples, while other models, like GANs and VAEs, only use decoder structures. Training costs for GANs and VAEs are similar to HUD because both require additional components like a discriminator or encoder. MPRNet, as a spectral super-resolution model, is cheaper since it directly generates HSIs from RGB inputs. Memory costs follow a similar pattern, with diffusion models requiring more memory due to their larger feature dimensions and iterative denoising process. For larger images (e.g., 256x256), computational complexity rises significantly, with 24 GB of memory insufficient for training models with batch size 4.

### v3

@@ Abstract
We address hyperspectral image (HSI) synthesis, a problem that has garnered growing interest yet remains constrained by the conditional generative paradigms that limit sample diversity. While diffusion models have emerged as a state-of-the-art solution for high-fidelity image generation, their direct extension from RGB to hyperspectral domains is challenged by the high spectral dimensionality and strict physical constraints inherent to HSIs. To overcome the challenges, we introduce a diffusion framework explicitly guided by hyperspectral unmixing. The approach integrates two collaborative components: (i) an unmixing autoencoder that projects generation from the image domain into a low-dimensional abundance manifold, thereby reducing computational burden while maintaining spectral fidelity; and (ii) an abundance diffusion process that enforces non-negativity and sum-to-one constraints, ensuring physical consistency of the synthesized data. We further propose two evaluation metrics tailored to hyperspectral characteristics. Comprehensive experiments, assessed with both conventional measures and the proposed metrics, demonstrate that our method produces HSIs with both high quality and diversity, advancing the state of the art in hyperspectral data generation.
@@ I Introduction
Hyperspectral image (HSI) generation represents a critical need in remote sensing. While HSIs enable a wide range of downstream applications [1, 2], their acquisition remains fundamentally limited by current sensor capabilities [3], resulting in a paucity of large-scale, high-fidelity datasets. This data scarcity continues to challenge methodological progress and robust evaluation in hyperspectral research [4]. In light of these limitations, the synthesis of diverse, realistic HSIs is emerging as a pivotal yet comparatively underexplored direction for advancing the field.
However, existing approaches are predominantly conditional and thus unsuitable for scalable data expansion [5]. These approaches typically employ conditional constraints, such as RGB images or segmentation maps, to guide the distribution of generated hyperspectral images [6, 7, 8]. While effective for spectral super-resolution, these methods fall short of generative modeling, as they prioritize spectral restoration over the comprehensive simulation of the underlying HSI distribution. While such conditioning can improve fidelity, it intrinsically curtails sample diversity and limits the generative capacity required for comprehensive data augmentation.
@@
To address these challenges, we integrate hyperspectral unmixing into the generative framework. We project HSIs into the abundance space, which is low-dimensional and governed by the explicit constraints, namely non-negativity and unity [15]. Operating in this physically grounded manifold allows a generator to synthesize HSIs that are both computationally tractable and physically consistent.
To implement the proposed framework, we introduce the Hyperspectral Unmixing Diffusion Model (HUD), which consists of two key components: an Unmixing AutoEncoder (UAE) pair and an Abundance Diffusion Module (ADM). To address the challenge of high dimensionality, we design the UAE module. UAE is initialized using a hyperspectral unmixing algorithm, where the abundances are derived either through rigorously constrained optimization equations or a linear approximation. Reconstruction is performed directly using the endmembers, ensuring efficient dimensionality reduction while preserving essential spectral information. To address the challenge of physical constraint, we construct the ADM module. The diffusion model iteratively generates samples on the abundance space. However, traditional diffusion models rely on Gaussian distributions to describe latent features, which inherently violate the unity and non-negativity constraints of the abundance space. To address this limitation, we shift the abundance space into an unconstrained domain, apply Gaussian-based sampling, and then map the generated samples back to the abundance space. This approach enables the generation of physically consistent and high-fidelity hyperspectral data while maintaining computational efficiency. Moreover, we introduce two novel metrics tailored for HSI generation: point fidelity and block diversity. Experimental results indicate that HUD consistently outperforms existing models on both proposed metrics and traditional metrics.
@@ II Related Work
We briefly review three relevant areas: diffusion-based generative modeling, HSI generation, and hyperspectral unmixing.
@@ II-A Diffusion Models
Denoising Diffusion Probabilistic Models (DDPMs) [16] have recently become central to state-of-the-art generative modeling. DDPMs define a forward noising process that progressively transforms data into standard Gaussian noise, and a learned reverse-time Markov process that maps noise back to data. The reverse transitions are parameterized by a denoising network, commonly a U-Net [17], trained to approximate the score or noise residual.
Latent Diffusion Models (LDMs) [18], popularized by Stable Diffusion, address the computational burden and limited resolution of pixel-space diffusion by operating in a compressed latent space. An autoencoder, typically pretrained on large-scale RGB datasets, projects images to a low-dimensional manifold where diffusion is performed, significantly improving efficiency and achievable resolution relative to pixel-space DDPMs.
@@ II-B HSI Generation
Recently, several studies have explored the use of generative models for HSI generation, including GAN, VAE, and Stable Diffusion [26, 27, 28, 29, 30, 31, 32, 33]. However, most of these approaches still rely on conditional images such as RGB or multi-spectral images during the generation process. This reliance limits the diversity of the generated hyperspectral images, as the spatial distribution is largely determined by the conditional input.
@@ III Method
Figure 1: Components of HUD.
In this section, we show the detailed design of the proposed model, including the unmixing autoencoder module and the abundance diffusion module. We first introduce the overall framework of the model, then describe the unmixing autoencoder module and the abundance diffusion module in detail. Finally, we introduce the evaluation metrics for HSI generation.
III-A Overall Framework
The proposed HUD consists of two key modules: the unmixing autoencoder module and the abundance diffusion module. The unmixing autoencoder module reduces the dimensionality of hyperspectral images by projecting them into a low-dimensional abundance space, while the abundance diffusion module generates samples corresponding with physical constraints.
Specifically, given a hyperspectral image dataset, we first extract the endmembers by a hyperspectral unmixing algorithm. Then, the endmembers are used to initialize the autoencoder for the diffusion module. To train the diffusion model, we first encode the hyperspectral images into the abundance space using the unmixing encoder, followed by a projection to relax the constraints of non-negativity and unity. The diffusion model is then trained to learn the distribution of the projected abundance maps. During sampling, we first sample from a standard Gaussian distribution, then apply the denoising process to generate projected abundance maps, and convert them back to the constrained abundance space. Finally, the decoder reconstructs the hyperspectral images from the abundance maps. The overall framework is illustrated in Figure˜1.
III-B Unmixing Autoencoder Module
The unmixing autoencoder can use any algorithm to extract endmember, and in this paper we use VCA for example. Abundances are subject to two key constraints: non-negativity and unity. Specifically, the abundance values for each pixel must be non-negative and sum to one, reflecting the physical interpretation that each pixel is composed of a mixture of endmembers in varying proportions. Therefore, given $Y$ and $A$, $X$ is usually solved through a constrained optimization problem of the following form:
[…]
@@ III-B Unmixing Autoencoder Module
III-C Abundance Diffusion Module
@@
where $Z$ represents the projected feature in an unconstrained space, and $\hat{X}$ is the reconstructed abundance map that satisfies the original constraints. This mapping is near-lossless for typical magnitudes:
@@
For simplification, we assume that $\Sigma_{\theta}(Z_{t},t)=\sigma^{2}_{t}\mathbf{I}$, and use reparameterizing $Z_{t}(Z_{0},\epsilon)=\sqrt{\bar{\alpha}_{t}}Z_{0}=\sqrt{1-\bar{\alpha}_{t}}\epsilon$ for $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, the middle term in Equation˜12 can be written as:
@@
[…]
@@
After training, we sample from standard Gaussian distribution to get $Z_{T}$ and follow the denoising process step by step. The training and sampling procedure are shown in Algorithms˜1 and 2.
@@
[…]
@@
III-D HSI Specialized Evaluation Metric
@@
[…]
@@ IV Experiments
In this section, we empirically demonstrate the superiority of HUD. We first introduce the datasets for the experiments, the comparison models, and the selection of hyperparameters. Then, we list images generated by different models. Finally, we analyze the performance of different models using quantitative metrics.
@@ IV-A Experiments Setup
We use Indianpines, KSC, Pavia, PaviaU, and Salinas to showcase the generation results. (1) Indianpines contains an image scene with spatial size $145\times 145$ and 220 bands, covering a wavelength range of [104108],[150163],220[104-108],[150-163],220 nm, having 16 classes. The Band 46,17,11 from the data are chosen for pseudo-color visualization. (2) KSC contains an image scene with spatial size $512\times 614$ and 176 bands, covering a wavelength range of 400 - 2500 nm, having 13 classes. The Band 28,9,10 from the data are chosen for pseudo-color visualization. (3) Pavia contains an image scene with spatial size $1093\times 715$ and 102 bands, having 9 classes. The Band 46,27,10 from the data are chosen for pseudo-color visualization. (4) PaviaU contains an image scene with spatial size $610\times 340$ and 103 bands, having 9 classes. The Band 46,27,10 from the data are chosen for pseudo-color visualization. (5) Salinas contains an image scene with spatial size $512\times 217$ and 204 bands, covering a wavelength range of [104108],[150163],220[104-108],[150-163],220 nm, having 16 classes. The Band 36,17,11 from the data are chosen for pseudo-color visualization.
HUD will be compared with the following models: VAE, GAN [44], MPRNet [45], UD [29] and UBF [46]. As there are few hyperspectral image generation algorithms, we adapted the classic models VAE and GAN from RGB generation tasks. Their first convolutional layer, usually a 1x1 conv layer, is resized according to the dimensions of the dataset. The other structures remain unchanged. We also evaluate the generation performance of MPRNet, a state-of-the-art hyperspectral image generation model, but it requires RGB images as conditions. Following the original paper, we directly input the pseudo RGB images to MPRNet. We also compare with UD and UBF, which are the state-of-the-art diffusion based generation models for hyperspectral images. The implementation of these models is based on the official code. We only change the dataloaders according to the datasets, and the structures and hyperparameters remain unchanged.
The UAE module is initialized separately for each dataset, and the dimensionality of $Z$ is determined by the number of categories in the dataset. The UAE is pretrained and does not participate in the training of the diffusion model. The time steps of the diffusion model are set to 1000. Except for the downsampling layer, the structures and hyperparameters of VAE, GAN, and MPRNet are set according to the original papers. All models are trained on a 4090 GPU. We randomly crop 32x32 overlapping sub-images from the original image as training samples.
@@ IV-B1 Results on Pseudo-Color Images
The pseudo-color visualization of 5 datasets is shown in Figure˜3. The first column shows the pseudo-color visualization of the real HSI, and the generated images from different methods are listed on the right side. The pseudo-color visualization can reflect the spatial distribution of HSIs. Specifically, the generated images from GAN have grids overlay on the surface, resembling a mosaic. This problem is probability caused by the downsampling layers in generators and discriminators, which overlook some specific areas because of the pooling or striding operation. VAE directly estimates the distribution of original images, but the scheme limits the performance of the model. Consequently, the generated images from VAE are blurry and lack details. As a spectral super-resolution model, MPRNet generates images with high quality, but the spatial distribution is limited to the original image. Although the autoencoders of UD and UBF are trained with physical constraint normalization, they cannot well restore the spectral information. As a result, UD and UBF have good spatial distributions but poor color visualization. HUD also generates high-quality images, and the spatial distribution is more diverse than MPRNet. For example, there are some new roads and buildings in the generated images of HUD, which are not present in the original images.
IV-B2 Results on Spectral Curves
The spectral curve distributions of the real HSI and the generated HSI are shown in Figure˜4. We stack all the spectral curves of an HSI to show the spectral distributions. The boundary shows whether the generated spectral curves are consistent with the real HSI, and the density (color) infers the distribution of endmembers. Generally, GAN, MPRNet and HUD generate spectral curves that are more consistent with the real HSI. GAN tends to generate spectral curves that are frequently seen in the real HSI to confuse the discriminator by high-quality samples, so the distribution of generated spectral curves are concentrated. VAE only estimates the main characteristics of the real HSI due to the deficient structure. MPRNet generates high-quality spectral curves, but the distribution is limited to the real HSI, so the generated images lack diversity. UD and UBF combines the physical constraints with reconstruction loss. The spectral distributions show similar patterns with real HSI, but there is a shift in the absolute value. HUD generates high-quality spectral curves that follows the overall shape and distribution of the real HSI, while there are also different distributions in the detail.
@@ IV-C1 Results on Fidelity and Diversity
$F_{p}$ $\uparrow$ $D_{b}$ $\downarrow$ $D_{b}/F_{p}$ $\downarrow$
Dataset GAN VAE MPRNet UD UBF HUD GAN VAE MPRNet UD UBF HUD GAN VAE MPRNet UD UBF HUD
Indianpines 0.996 0.914 0.996 0.558 0.559 0.999 0.981 0.914 0.995 0.553 0.553 0.978 0.984 1.000 0.999 0.991 0.989 0.979
KSC 0.998 0.978 0.983 0.836 0.831 0.999 0.988 0.978 0.983 0.833 0.830 0.990 0.989 1.000 1.000 0.997 0.999 0.991
Pavia 0.996 0.941 0.999 0.867 0.836 0.999 0.949 0.936 0.942 0.818 0.795 0.912 0.953 0.995 0.943 0.943 0.951 0.913
PaviaU 0.996 0.941 0.997 0.875 0.866 0.999 0.948 0.934 0.984 0.823 0.820 0.919 0.952 0.992 0.987 0.941 0.947 0.920
Salinas 0.995 0.998 0.998 0.553 0.619 0.999 0.984 0.982 0.988 0.545 0.610 0.981 0.989 0.984 0.990 0.986 0.985 0.982
Avg 0.994 0.954 0.995 0.537 0.742 0.999 0.968 0.949 0.978 0.714 0.722 0.958 0.969 0.994 0.984 0.972 0.974 0.959
@@
Metric Dataset GAN VAE MPRNet UD UBF HUD
IS $\uparrow$ Indianpines 1.08 1.02 1.04 1.00 1.00 1.07
KSC 1.00 1.00 1.00 1.00 1.00 1.00
Pavia 1.01 1.01 1.04 1.00 1.00 1.01
PaviaU 1.03 1.03 1.04 1.00 1.00 1.05
Salinas 1.05 1.04 1.00 1.00 1.01 1.05
Avg 1.03 1.02 1.03 1.00 1.00 1.04
@@
Metric Dataset GAN VAE MPRNet UD UBF HUD
FID $\downarrow$ Indianpines 8.03 240.91 20.98 124.41 171.62 4.71
KSC 10.87 5.45 4.76 51.99 53.68 6.69
Pavia 4.81 13.82 2.37 23.37 19.01 4.79
PaviaU 4.92 9.13 3.13 14.46 11.32 2.81
Salinas 15.03 21.33 29.09 76.63 91.64 23.12
Avg 8.73 58.13 12.07 58.17 69.45 8.42
The quantitative results are shown in Tables˜II, III and I. According to Table˜II, all models achieve similar results while HUD performs slightly better than the other models. According to Table˜III, the results are unstable across different datasets and models. IS and FID are originally proposed for RGB images, which requires a large-scale pretrained InceptionV3. However, the hyperspectral images have significant differences in distribution, so the classification networks should be trained on each dataset. As a result, the metrics are unstable and questionable. Regardless, HUD still achieves the top results in the IS and FID comparison.
As shown in Table˜I, HUD achieves the highest point fidelity, indicating that its generated images better preserve the spectral content of real images. In contrast, although UD and UBF introduce the physical constraints into data-driven autoencoders, they cannot generate high quality spectral pixels. This highlights the effectiveness of directly applying unmixing algorithm in HSI generation. MPRNet, a spectral super-resolution model, ranks second in point fidelity. UD, UBF and VAE generates the most diverse images, but their poor point fidelity result in overall low-quality generations. HUD achieves the better block diversity compared with GAN and MPRNet, demonstrating richer spatial distributions. MPRNet ranks lowest in block diversity, suggesting its generated images are overly constrained by the original spatial patterns. Combining both point and block diversity, HUD delivers the best overall performance, followed by GAN. This indicates that HUD not only generates high-quality images but also maintains a diverse distribution.